1. An individual has a cardiac output of 5.25 L / min. The diameter of their aorta is estimated to be 22 mm and the cross-sectional area of the vena cavae is 12.5 cm2. What is the velocity of blood flow in the aorta relative to the velocity of blood flow in the vena cavae? Show your work.

2. An individual has a hepatic blood flow of 750 mL / min. The hepatic arterial pressure is 95 mm Hg and the hepatic venous pressure is 10 mm Hg. What is the vascular resistance of the liver for this individual? Show your work.

3. In the skeletal muscles of an individual, the following Starling pressures are measured: Pc = 35 mm Hg, Pi = 2 mm Hg, pc = 28 mm Hg, pc = 4 mm Hg. If the Kf is 0.45 mL / min · mm Hg, what is the direction and magnitude of fluid movement across the capillary? Show your work.

I have been struggling with these calculations. Any help is greatly appreciated so that I can understand how to do them. Thank you so much!!
Answer #1

1.flow = velocity ×area

Since flow is constant

Velocity is inversely proportional to area

That means velocity is inversely proportional to area(pie r²)

Velocity in aorta /venacava

= area of venacava ²/aorta²

= 12.5/15 (area = pie r²)

=5/6

Velocity of venacava /aorta= 6/5

= 1.2 times